Find equations of all lines having slope -1 that are tangent to the curve 𝑦=1/𝑥−1



1
Expert's answer
2021-04-27T01:24:26-0400

Given the equation "y=\frac{1}{x-1},"


"m=\dfrac{dy}{dx}=-\frac{1}{(x-1)^2}=-1 \qquad \text{(given)}"

"\implies \frac{1}{(x-1)^2} = 1\\
1= (x-1)^2\\
1= x^2-2x+1\\
1-1=x^2-2x\\
0=x(x-2)\\
\implies x=0 \text{ and } x=2"

For "x=0,"

"y=\frac{1}{0-1}=\frac{1}{-1}=-1\\
\therefore \text{the tangent line touches the curve at the point (0,-1)}"


The equation of the line is:

"y-y_1=m(x-x_1)\\
y-(-1)=-1(x-0)\\
y+1=-x\\
\implies x+y=-1 \qquad \cdots (i)"

at "x=2,"

"y=\frac{1}{2-1}=\frac{1}{1}=1\\
\therefore \text{the tangent line touches the curve at the point (2,1)}"

"y-y_1=m(x-x_1)\\
y-(-1)=-1(x-2)\\
y-1=-x+2\\
\implies x+y=3 \qquad \cdots (ii)"

Thus, eqn (i) and (ii) are the required equations of the line that are tangential to the curve.